# Itô’s formula for flows of conditional measures on semimartingales

### v2

@@ 3.4. Properties of the conditional independent copies of semimartingales
$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}},\text{ for }t\in[0,T],$
therefore, by (2c) of Theorem 3.3, ${\mathrm{d}}\bm{Z}_{s}={\mathrm{d}}\big{(}g^{(1)}(\bm{X}_{s}),\cdots,g^{(n)}(\bm{X}_{s})\big{)}^{\mathcal{G}})$, we have
@@
$+\sum_{0<s\leq t}\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}}\bigg{|}{\mathcal{F}}\Bigg{]}$
For $J_{2}$, computing $[g^{(i)}(\bm{X}^{\prime}),g^{(j)}(\bm{X}^{\prime\prime})]_{t}$ for $1\leq i,j\leq n$, we have
[…]
@@
and computing $[g^{(i)}(\bm{X}^{\prime}),Y^{(k)}]_{t}$ for $1\leq i\leq n$, $1\leq k\leq d$, we have
[…]
@@
Consequently, using (2d) of Theorem 3.3, we have
@@
Now, to deal with $J_{3}$, note that
@@
where the third equation holds by the conditional dominated convergence theorem. Therefore, we can rewrite $J_{3}$ as
@@
since $\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $Y^{(k)}_{s}$ and $Y^{(k)}_{s-}$ are ${\mathcal{F}}$-measurable random variables for $s\in(0,t]$, $1\leq i,j\leq n$ and $1\leq k\leq d$. Recall that $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}$, simple calculation of linear derivatives gives us a version of the derivatives by
@@
Using the argument at the end of [24, Section 3.2.1], one can cancel out redundant terms in $J_{2}$ and $J_{3}$ to get the first summation in the right hand side of (2.3). Therefore, combining (4.6), (4.7), and (4.8) and plugging them back into (4.2) complete the proof.
@@
We claim that for a sequence of ${\mathbb{F}}$ stopping time $\{\tau_{n}\}_{n\in{\mathbb{N}}}$ such that $\tau_{n}\to T$, in order to prove that (2.3) holds for $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ with the process $(\mu,\bm{Y})=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$, it suffices to prove that (2.3) holds for $\Phi$ with the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$.
Indeed, suppose (2.3) holds for the truncated process, for $t<T$. Let $n\to\infty$, it is not hard to see that the left hand side of (2.3) converges and last three terms of the right hand side of (2.3) converge to the desired equation. Now, let us split the conditional expectation term of the right hand side of (2.3) into several terms. First, in order to prove that
@@
almost surely as $n\to\infty$, we separate it by,
@@
$\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{s}-\sum_{0<s\leq t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot\Delta\bm{X}^{\prime}_{s}$
@@
where $p^{\prime}=\frac{p}{p-1}$, $c_{1},c_{p}$ are constants in Proposition 2.2 and $C$ is a generic constant that may vary line by line. The first inequality is due to Proposition 2.2, the second inequality holds because of Proposition 2.3, the third inequality holds by $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, and the last one holds because $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Meanwhile, the second term of the right hand side of (4.10) is bounded by
@@
where $C$ is a generic constant that may vary line by line, and the right hand side of the above inequality is integrable due to the fact that $\|\bm{X}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{X}\|_{{\mathcal{H}}_{p}}<\infty$, $\|\bm{Y}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{Y}\|_{{\mathcal{H}}_{p}}<\infty$ and the assumption on $\bm{X}$. Therefore, using the conditional dominated theorem, we conclude (4.9). Next, to prove that
@@
almost surely as $n\to\infty$, we note that
@@
$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{u}\bigg{|}\leq\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}$
@@
where $c_{p^{\prime\prime}}$ is the constant in Proposition 2.2, $C$ is a generic constant that may vary from line to line and $p^{\prime\prime}=\frac{p-2}{2p}$. Thus, (4.12) holds due to the conditional dominated theorem, we conclude . Similar arguments holds also for the convergence of $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]_{s}$ and $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]_{s}$.
@@
With simple application of the fundamental theorem of calculus for the linear derivative, we have
@@
where $C$ is a generic constant that may vary from line by line. Therefore, the right hand side of the above inequality is integrable and the convergence of the conditional expectation of (4.13) is ensured by conditional dominated convergence theorem. Similar arguments can be applied to deduce the convergence of the conditional expectation of $\big{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\big{)}\cdot\Delta\bm{Y}_{s}$ and $\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$, s and thus we conclude that (2.3) holds for the original process $(\mu,\bm{Y})$ for $0\leq t<T$.
For the case of $t=T$, taking $n\to\infty$, left hand side equals $\Phi(\mu_{T-},\bm{Y}_{T-})-\Phi(\mu_{0},\bm{Y}_{0})$ while all terms at the right hand sides is evaluated at $T$ except the last term being
@@
The following result slightly generalizes that in [16, Theorem 4.4, Theorem 4.10, Section 4] as it includes the convergences of the derivatives of the linear derivative.
@@
Let $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, then there exist a sequence of $C^{2,2}$ cylinder functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$ such that one has the point-wise convergence
@@
as $n\to\infty$, for all $\bm{x}_{1},\bm{x}_{2},\bm{y},\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. Moreover, there exists a constant $c>0$ such that
@@
$\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c_{K}\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{p-2}\big{)},\qquad\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{{p-2}}+|\bm{y}|^{p-2}\big{)},$
@@
The proof is similar to that of [16, Section 4]. The main step of their proof consists of constructing an operator $T_{n}:C_{b}({\mathbb{R}}^{d})\to C_{b}({\mathbb{R}}^{d})$ and the corresponding adjoint $T_{n}^{*}:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\to{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ corresponding square $T_{n}^{\otimes 2}:C_{b}({\mathbb{R}}^{2d})\to C_{b}({\mathbb{R}}^{2d})$ such that $T_{n}\varphi_{n}\to\varphi$, when $\varphi_{n}\to\varphi$ and $T_{n}^{*}\mu\to\mu$. Define $f_{n}(\mu,\bm{y}):=\Phi(T_{n}^{*}\mu,\bm{y})$, then the key identities to be established are
$\frac{\delta f_{n}}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=T_{n}\Big{(}\frac{\delta\Phi}{\delta\mu}(T_{n}^{*}\mu,\bm{y},\cdot)\Big{)}(\bm{x}_{1}),\quad\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=T_{n}^{\otimes 2}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(T_{n}^{*}\mu,\bm{y},\cdot,\cdot)(\bm{x}_{1},\bm{x}_{2}).$
for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$ and $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. To this end, it suffices to note that the construction of $T_{n}$ ensures
@@
for all $\varphi\in C^{1}_{b}({\mathbb{R}}^{d})$ and $\psi\in C^{2}_{b}({\mathbb{R}}^{2d}).$

### v3

@@ 3.5. Properties of the conditionally independent copies of semimartingales
$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}},\text{ for }t\in[0,T].$
Therefore, by (1) of Theorem 3.5, ${\mathrm{d}}\bm{Z}_{s}={\mathrm{d}}\big{(}{\mathbb{E}}[g^{(1)}(\bm{X}_{s})|{\mathcal{G}}],\cdots,{\mathbb{E}}[g^{(n)}(\bm{X}_{s})|{\mathcal{G}}]\big{)}$, and
@@
$+\sum_{0<s\leq t}\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}}\bigg{|}{\mathcal{F}}\Bigg{]}.$
For $J_{2}$, we see that
$\big{[}$ $g^{(i)}(\bm{X}^{\prime}_{t}),g^{(j)}(\bm{X}^{\prime\prime}_{t})\big{]}_{t}=\big{[}(g^{(i)}(\bm{X}^{\prime}))^{c},(g^{j}(\bm{X}^{\prime\prime}))^{c}\big{]}_{t}+\sum_{0<s\leq t}\Delta g^{(i)}(\bm{X}^{\prime}_{s})\Delta g^{(j)}(\bm{X}^{\prime\prime}_{s})$
@@
and
$\big{[}$ $g^{(i)}(\bm{X}^{\prime}_{t}),Y^{(j)}\big{]}_{t}=\big{[}(g^{(i)}(\bm{X}^{\prime}))^{c},(Y^{j})^{c}\big{]}_{t}+\sum_{0<s\leq t}\Delta g^{(i)}(\bm{X}^{\prime}_{s})\Delta Y^{(j)}_{s}$
@@
Consequently, by (2) of Theorem 3.5,
@@
To deal with $J_{3}$, note that
@@
where the third equality holds by the conditional dominated convergence theorem (Lemma 2.5). Therefore, we can rewrite $J_{3}$ as
@@
since $\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $Y^{(k)}_{s}$ and $Y^{(k)}_{s-}$ are $\mathcal{F}$}-measurable random variables for $s\in(0,t]$, $1\leq i,j\leq n$, and $1\leq k\leq d$.
Recall that $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}$, simple calculation of linear derivatives yields a version of the derivatives by
@@
Using the argument at the end of [31, Section 3.2.1], one can cancel out redundant terms in $J_{2}$ and $J_{3}$ to get the first summation in the right hand side of (2.3). Therefore, combining (4.6), (4.7), and (4.8) and plugging them back into (4.2) completes the proof.
@@
Take a sequence of $\mathbb{F}$}-stopping time $\{\tau_{n}\}_{n\in{\mathbb{N}}}$ such that $\tau_{n}\to T$, in order to prove (2.3) for $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ with the process $(\mu,\bm{Y})=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$, it suffices to prove (2.3) for $\Phi$ with the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$. Indeed, suppose (2.3) holds for the truncated process, for $t<T$. Let $n\to\infty$, it is not hard to see that the left hand side of (2.3) converges and the last three terms of the right hand side of (2.3) converge to the desired equation.
Now, let us split the conditional expectation term of the right hand side of (2.3) into several terms. First, in order to prove that
@@
a.s. as $n\to\infty$, we separate it by,
@@
$\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{s}-\sum_{0<s\leq t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot\Delta\bm{X}^{\prime}_{s}.$
@@
where $p^{\prime}=\frac{p}{p-1}$, $c_{1}$ and $c_{p}$ are constants in Proposition 2.2, and $C$ is a generic constant that may vary line by line. The first inequality is due to Proposition 2.2, the second inequality is by Proposition 2.3, the third inequality is from $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, and the last one holds because $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Meanwhile, the second term of the right hand side of (4.10) is bounded by
@@
where $C$ is a generic constant, and the right hand side of the above inequality is integrable because $\|\bm{X}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{X}\|_{{\mathcal{H}}_{p}}<\infty$, $\|\bm{Y}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{Y}\|_{{\mathcal{H}}_{p}}<\infty$, and by the assumption of $\bm{X}$. Therefore, by the conditional dominated convergence theorem (see Lemma 2.5), we conclude (4.9).
Next, to prove that
@@
a.s. as $n\to\infty$, note that
@@
$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{u}\bigg{|}\leq\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T},$
@@
where $c_{p^{\prime\prime}}$ is the constant in Proposition 2.2, $C$ is a generic constant that may vary from line to line and $p^{\prime\prime}=\frac{p-2}{2p}$. Thus, (4.12) holds by Lemma 2.5. Similar arguments holds also for the convergence of $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]_{s}$ and $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]_{s}$.
@@
Applying the fundamental theorem of calculus for the linear derivative, we have
@@
where $C$ is a generic constant. The right hand side of the above inequality is integrable and the convergence of the conditional expectation of (4.13) is ensured by Lemma 2.5. Similar arguments can be applied to deduce the convergence of the conditional expectation of $\big{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\big{)}\cdot\Delta\bm{Y}_{s}$ and $\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$, and thus we conclude that (2.3) holds for the original process $(\mu,\bm{Y})$ for $0\leq t<T$.
For the case of $t=T$, taking $n\to\infty$, then the left hand side equals $\Phi(\mu_{T-},\bm{Y}_{T-})-\Phi(\mu_{0},\bm{Y}_{0})$ while all terms on the right hand side are evaluated at $T$ except for the last term which is
@@
The following result slightly generalizes that in [21, Theorem 4.4, Theorem 4.10, Section 4] as it also includes the convergences of the associated linear derivatives.
@@
Let $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, then there exists a sequence of $C^{2,2}$ cylinder functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$ such that one has the point-wise convergence
@@
as $n\to\infty$, for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. Moreover, there exists a constant $c>0$ such that
@@
$\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{p-2}\big{)},\qquad\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{{p-2}}+|\bm{y}|^{p-2}\big{)},$
@@
The proof is similar to that of [21, Section 4]. Following their construction procedure, fix $n\in{\mathbb{N}}$ and cover the ball $B_{n}:=\{\bm{x}\in{\mathbb{R}}^{d}:|\bm{x}|\leq n\}$ by finitely many open sets of diameter at most $1/n$, denoted by $U^{n}_{i}$, with $i=1,...,N_{n}$. Set $U^{n}_{0}={\mathbb{R}}^{d}\setminus B_{n}$ to get an open cover in ${\mathbb{R}}^{d}$. Then, fix points ${\bm{x}}^{n}_{i}$ in $\overline{U^{n}_{i}}$ with the minimal norm for $i=0,1,..,N_{n}$. Next, define a family of functions $\{\psi^{n}_{i}\}$ as a partition of unity subordinate to $\{U^{n}_{i}\}$. (That is, each $\psi^{n}_{i}$ is a continuous function supported on $U^{n}_{i}$, such that $\sum_{i=0}^{N_{n}}\psi^{n}_{i}(\bm{x})=1$ for all $\bm{x}\in{\mathbb{R}}^{d}$). Finally, define the operator $T_{n}:C_{b}({\mathbb{R}}^{d})\to C_{b}({\mathbb{R}}^{d})$ such that for any function $\varphi$ on ${\mathbb{R}}^{d}$,
$T_{n}\varphi(x)=\sum_{i=0}^{N_{n}}\varphi({\bm{x}}^{n}_{i})\psi^{n}_{i}(\bm{x}).$
Its adjoint operator $T_{n}^{*}:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\to{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ and the tensor square operator $T_{n}^{\otimes 2}:C_{b}({\mathbb{R}}^{2d})\to C_{b}({\mathbb{R}}^{2d})$ are defined accordingly. Now, define $f_{n}(\mu,\bm{y}):=\Phi(T_{n}^{*}\mu,\bm{y})$, then $f_{n}$ are $C^{2,2}$ cylindrical functions and the key identities to be established are
$\frac{\delta f_{n}}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=T_{n}\Big{(}\frac{\delta\Phi}{\delta\mu}(T_{n}^{*}\mu,\bm{y},\cdot)\Big{)}(\bm{x}_{1}),\quad\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=T_{n}^{\otimes 2}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(T_{n}^{*}\mu,\bm{y},\cdot,\cdot)(\bm{x}_{1},\bm{x}_{2}),$
for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$ and $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ (see [21, Lemma 4.6, Lemma 4.12]). To this end, note that the construction of $T_{n}$ ensures
@@
for all $\varphi\in C^{1}_{b}({\mathbb{R}}^{d})$ and $\psi\in C^{2}_{b}({\mathbb{R}}^{2d}).$ Thus repeating the proof of [21, Theorem 4.4, Theorem 4.10] finishes the proof of the proposition.